OpenStudy (anonymous):

Solve for x: |3x + 2| = 2 please explain it to me!

OpenStudy (anonymous):

Since you have the expression 3x + 2 within absolute value signs, that expression will equal both 2 and -2, so you will have 2 equations. 3x + 2 = 2 and 3x + 2 = -2 So, x = 0 and -4/3

OpenStudy (anonymous):

That is because |-2| = |2| = 2

OpenStudy (anonymous):

ok thank you soo much! and on this one. Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 am i right?

OpenStudy (anonymous):

Yes! Very good!

OpenStudy (anonymous):

am sorry last one! Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 am not sure which ones right btw thanks for all the help

OpenStudy (anonymous):

|x| = 8 - 5 = 3, so, x = 3 and -3